OpenStudy (anonymous):

A fish tank is half full of water. if ten gallons of water were added, the tank will be 7/8 full. What is the capacity of the of the tank in gallons?

OpenStudy (turingtest):

(1/2)x+10=(7/8)x 4x+80=7x 3x=80 x=80/3 gallons

OpenStudy (turingtest):

where x is the capacity of the tank

OpenStudy (anonymous):

initially the tank was half full. on adding ten gallons the tank was 7/8 full so ten gallons equals 7/8-1/2 of the tank = 3/8 hence full tank would have the capacity of 10*8/3=80/3 gallons